OpenStudy (anonymous):

If f(x) = 3x + 5 and g(x) = 2x - 9, find f(x) - g(x).

OpenStudy (anonymous):

just do the abstraction

OpenStudy (anonymous):

f(x) = 3x + 5, g(x) = 2x - 9 f(x)-g(x)=___________fill this

OpenStudy (bradely):

f(x)=3x+5 g(x)=2x-9 f(x)-g(x)=(3x+5)-(2x-9) =3x-2x+5+9 =x+14 Source: http://www.mathskey.com/question2answer/